OpenStudy (anonymous):

Solve x^2 + 6x + 7 = 0. quadratics using the quadratic formula

OpenStudy (vijay):

root2 -3

OpenStudy (vijay):

& -3 -root2

OpenStudy (anonymous):

possible answers are x=-1andx=-5 , 3+_sqrt2, -3+-sqrt2, -3+-sqrt2/2

OpenStudy (vijay):

-1 and -5 not possible

OpenStudy (anonymous):

Using the quadratic equation: ax^2 + bx + c = 0 we have: a =1 b = 6 c = 7 Then we plug into the quadratic formula: \[-b \pm \sqrt{b^{2}-4ac}/2a\] We get the following solutions: x = -1, -5, 3 + sqrt(2), 3 -sqrt(2), Hope this helps. If it does, please choose good answer. Visit my website at http://www.tutorsean.net for free math training videos.